# DD-VNB: A Depth-based Dual-Loop Framework for Real-time Visually Navigated Bronchoscopy

### 04 Mar 2024 v1

@@ I Introduction
To achieve generalization across cases, we introduce a novel depth estimation network. Initially, the depth estimation maps endoscope frames to the depth space, serving as input for the pose inference module. This process eliminates patient-specific textures, ensuring the generalization ability of the localization method. To address the challenge of estimating bronchoscopic depth within specific scale, we propose a knowledge-embedded depth estimation network. We innovatively incorporates view synthesis into an unpaired image translation framework, leveraging the geometric principles of view synthesis to inform and constrain the learning process. This approach is motivated by earlier research that utilized view synthesis for depth and ego-motion estimation [15], as well as recent efforts to apply geometry view synthesis in creating realistic surgical scenes [16].
@@
    A fast localization module is proposed that embeds a efficient ego-motion estimation network within the loop of single-view depth map registration, enabling fast and accurate pose inference.
@@ III Method
Fig. 2 shows an overview of the proposed bronchoscopic localization framework. The localization methodology involves depth estimation and localization. Learning-based depth estimation is accomplished based on a adversarial architecture with knowledge embedding as supervision. The localization part consists two algorithms: a ego-motion estimation network that takes two nearby bronchoscopic depth maps as input to infer a 6 DOF relative pose between them; and a depth map registration method mapping a single estimated depth map to the pre-operative airway model, correcting the accumulative error of incremental ego-motion inference.
@@ III-A Depth Estimation with View Synthesis as Supervision
where $T_{t-n,t}$ and $\boldsymbol{R}_{t-n,t}$ are the translation vector and rotation matrix from $t-n$ to $t$. Note that $n$ should be carefully chosen for sufficient co-visibility between the two time points.
@@
III-B Depth Based Ego-motion Inference
As depth map and corresponding camera poses are accessible in virtual bronchoscopy, the learning objective is to minimize the difference between predicted transformation $\widehat{P}_{t-n,t}$ and ground truth transformation $P_{t-n,t}$. Therefore, the relative camera pose inference loss is defined as:
@@ III-B Depth Based Ego-motion Inference
Following [29, 30], FlownetC encoder [31] is utilized to extract image features. Then, we use five convolutional blocks to regress a pose vector from extracted features. By using virtual depth maps and poses as training data, our ego-motion network can be deployed directly to estimated depth map of incoming bronchoscopic frames during test time.
III-C Depth Map Registration
Relying on ego-motion estimation network alone for tracking yields accumulative error, as errors of previous inferences will propagate over time to current estimation in incremental localization. Thus, refining previous results becomes necessary. Registration between estimated depth maps and pre-operative bronchi model provides absolute position of bronchoscope, eliminating accumulative error from previous relative pose inference.
@@ VI Conclusion
Our study introduces a bronchoscopic localization framework, featuring a knowledge-embedded depth estimation network within a dual-loop scheme for fast and accurate localization. Utilizing monocular frames for depth map inference and integrating both ego-motion estimation and airway CT registration, our work marks a stride toward real-time, learning-based bronchoscopic localization capable of adapting to unseen airways, with future enhancements aimed at further accuracy improvements through feature matching and relocalization strategies.

### 15 Mar 2024 v2

@@ Abstract
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
To achieve generalization across cases, we propose a knowledge-embeded depth estimation network. Initially, the depth estimation maps endoscope frames to the depth space, serving as input for the pose inference module. This process eliminates patient-specific textures, ensuring the generalization ability of the localization method. Moreover, we innovatively incorporates view synthesis into an unpaired image translation framework, leveraging the geometric principles of view synthesis to inform and constrain the learning process. This approach is motivated by earlier research that utilized view synthesis for joint training of unsupervised depth and motion estimation [15], as well as recent efforts to apply geometry view synthesis in creating realistic surgical scenes [16].
@@
    A fast localization module is proposed that embeds an efficient ego-motion estimation network within the loop of single-view depth map registration, enabling fast and accurate pose inference.
@@ III Method
Fig. 2 shows an overview of the proposed bronchoscopic localization framework. The localization methodology involves depth estimation and localization. Learning-based depth estimation is accomplished based on a adversarial architecture with knowledge embedding as supervision. The localization part consists two algorithms: an ego-motion estimation network that takes two nearby bronchoscopic depth maps as input to infer a 6 DOF relative pose between them; and a depth map registration method mapping a single estimated depth map to the pre-operative airway model, correcting the accumulative error of incremental ego-motion inference.
@@ III-A Depth Estimation with View Synthesis as Supervision
where $T_{t-n,t}$ and $\boldsymbol{R}_{t-n,t}$ are the translation vector and rotation matrix from $t-n$ to $t$.
@@
III-B Dual-loop Localization
Our localization module integrates an ego-motion estimation network into the loop of depth registration for real-time bronchoscope tracking, enhancing accuracy by using relative pose changes as initial guesses in depth registration against the pre-operative 3D model.
Depth Based Ego-motion Inference. As depth map and corresponding camera poses are accessible in virtual bronchoscopy, the learning objective is to minimize the difference between predicted transformation $\widehat{P}_{t-n,t}$ and ground truth transformation $P_{t-n,t}$. Therefore, the relative camera pose inference loss is defined as:
@@ III-B Dual-loop Localization
We use FlownetC encoder [29] to extract image features. Then, five convolutional blocks are utilized to regress a pose vector from extracted features. By using virtual depth maps and poses as training data, our ego-motion network can be deployed directly to estimated depth map of incoming bronchoscopic frames during test time.
Depth Map Registration. Relying on ego-motion estimation network alone for tracking yields accumulative error, as errors of previous inferences will propagate over time to current estimation in incremental localization. Thus, refining previous results becomes necessary. Registration between estimated depth maps and pre-operative bronchi model provides absolute position of bronchoscope, eliminating accumulative error from previous relative pose inference.
@@ VI Conclusion
Our study proposes a bronchoscopic localization framework, featuring a knowledge-embedded depth estimation network within a dual-loop scheme for fast and accurate localization. Utilizing monocular frames for depth map inference and integrating both ego-motion estimation and airway CT registration, our work marks a stride toward real-time, learning-based bronchoscopic localization capable of adapting to unseen airways, with future enhancements aimed at further accuracy improvements through feature matching and relocalization strategies.